OpenStudy (anonymous):

Is each ordered pair a solution of 4x – 3y . 6? Explain. (0,2)

OpenStudy (johnweldon1993):

What you have is an ordered pair (0 , 2) (x , y) This means that 0 = x 2 = y So what we do...is we take our equation 4x - 3y = 6 And everywhere we see an 'x' we replace it with 0 ...and everywhere we see a 'y' we replace it with a 2 So that becomes 4(0) - 3(2) = 6 0 - 6 = 6 -6 = 6 That is not true since \(\large -6 \cancel{=} 6\) So this point is NOT a solution to the equation

OpenStudy (anonymous):

thanks for your help

OpenStudy (johnweldon1993):

No problem